OpenStudy (anonymous):

which of the binomials below is a factor of this trinomial? x^2+4x-60 A)x+6 B)x^2-6 C)x-6 D)x-10

OpenStudy (akashdeepdeb):

So watcha waiting for? factorize!! :D x2 + 4x - 60 So write 4x in the form of factors of 60 Like this 10x-6x right? [BECAUSE 10*6 = 60] Getting me? So x2 + 10x - 6x - 60 Group first, second And third fourth to get The 2 factors!! :D

OpenStudy (akashdeepdeb):

If the Quadratic expression is Ax2 + Bx + c [WHERE ALPHA AND BETA ARE ROOTS] Then alpha + beta should be -B/A And alpha*beta = C/A Getting me? :)